OpenStudy (anonymous):

Easy equation problem Just need to make an equation and solve for this. It takes Joey 4 hours to finish refurbishing a room and it takes Brynette 6 hours. If they were to work together, how would it take to refurbish the room?

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 } + \frac{ 1 }{ 6 } = \frac{ 1 }{ x }\] I don't know if that equation would work but the 1/4 reps Joeys amount he could finih and 1/6 reps brynettes portion. x represents the number of hours it would take them to finish it together

OpenStudy (ranga):

Yes, you are on the right track. Solve for x.

OpenStudy (anonymous):

how would i do that?

OpenStudy (ranga):

Add the two fractions on the left. Then flip both sides to find x.

OpenStudy (anonymous):

do i cross multiply it to equal 10x or \[\frac{ 10 }{ x }\] ?

OpenStudy (the_fizicx99):

Need them to have the same denominators

OpenStudy (ranga):

1/4 + 1/6 = 6/24 + 4/24 = (6+4) / 24 = 10 / 24 = 5 /12 = 1 / x x = 12/5 = 2.4 hours.

OpenStudy (ranga):

LCM is 12 but I just cross multiplied.

OpenStudy (the_fizicx99):